# Learning Contrastive Feature Representations for Facial Action Unit Detection

### v6

@@ 5 Conclusion
In this paper, we propose a discriminative contrastive learning framework tailored for AU detection, introducing a novel contrastive loss named AUNCE. Building upon the InfoNCE loss, AUNCE refines the task by shifting focus from learning pixel-level information of the entire face to capturing subtle feature variations during AU activation, thus improving detection efficiency. We also introduce a positive sample sampling strategy that constructs three types of positive sample pairs, selectively augmenting positive samples, especially in the presence of noisy and false AU labels. This approach ensures the model is trained on more reliable examples, improving upon conventional methods that lack adaptive sampling. To address class imbalance in AU detection, we implement a negative sample re-weighting strategy that prioritizes minority AUs, helping the model focus on hard-to-detect samples. This strategy promotes hard example mining by giving more weight to underrepresented classes, enhancing model performance and generalization. Experimental results demonstrate that AUNCE consistently outperforms existing methods, particularly those utilizing weighted cross-entropy loss.
However, hyperparameter tuning remains a limitation of our study. While the proposed method demonstrates promising results, we found that performance is highly sensitive to the values of certain hyperparameters, with optimal settings varying across different datasets. The parameter tuning process is largely driven by empirical experience gained from extensive experimentation, which results in a tuning scheme that lacks the necessary flexibility. To address this limitation, our future work will focus on the development of adaptive hyperparameter selection techniques that can dynamically adjust to different datasets and conditions, enabling more robust and generalized performance across a variety of tasks and settings.
@@ Acknowledgement
During the preparation of this work the author(s) used $chatgpt$ in order to polish our paper. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.

### v7

@@ 4.4.1 $p_{1},p_{2},p_{3},p_{4}$, the probabilities of various positive sample pairs
The experimental results are presented in Table 11. When $p_{1}$=0.15, $p_{2}$=0.15, $p_{3}$=0.7, and $p_{4}$=0, the experimental result is optimal. The InfoMin [40] principle provides a theoretical explanation for the above results. The core of the InfoMin principle is to provide a criterion for judging the quality of a positive sample: a high-quality positive sample should retain the essential information required for the downstream task (AU detection) while minimizing its mutual information with the anchor as much as possible. Here, the ”anchor” refers to the original image or the core features of the target AU, and the ”positive examples” correspond to the samples represented by $p_{1}$, $p_{2}$, and $p_{3}$. For $p_{4}$-related samples, their mutual information with the anchor is extremely low due to their intrinsic noise nature, thus explaining why $p_{4}=0$ in the optimal configuration. The $p_{3}$-related samples (mixed positive samples) have moderate mutual information with the anchor, which balances the retention of positive class common features and the avoidance of information redundancy. As for $p_{1}$- and $p_{2}$-related samples, although they can filter noise, their mutual information with the anchor usually exceeds the ”sweet spot”. An increase in their proportion will lead to information redundancy and subsequent performance degradation, which justifies their optimal value of 0.15.
@@ 5 Conclusion
In this paper, we propose a discriminative contrastive learning framework tailored for AU detection, introducing a novel contrastive loss named AUNCE. Building upon the InfoNCE loss, AUNCE refines the task by shifting focus from learning pixel-level information of the entire face to capturing subtle feature variations during AU activation, thus improving detection efficiency. We also introduce a positive sample sampling strategy that constructs three types of positive sample pairs, selectively augmenting positive samples, especially in the presence of noisy and false AU labels. This approach ensures the model is trained on more reliable examples, improving upon conventional methods that lack adaptive sampling. To address class imbalance in AU detection, we implement a negative sample re-weighting strategy that prioritizes minority AUs, helping the model focus on hard-to-detect samples. This strategy promotes hard example mining by giving more weight to underrepresented classes, enhancing model performance and generalization. Experimental results demonstrate that AUNCE consistently outperforms existing methods, particularly those utilizing weighted cross-entropy loss. While current validation is limited to static-image datasets, it is worth noting that these datasets are derived from dynamic video sources. Hence, the proposed framework can be naturally extended to dynamic AU sequences.
Despite the proven effectiveness of our approach, hyperparameter selection emerges as a notable limitation. Model performance is sensitive to specific parameters, with optimal configurations varying across datasets—particularly the mixing ratio of positive samples. Its optimal value fluctuates significantly with changes in data quality and noise rates in real-world scenarios, yet the current determination of this ratio still relies on empirical adjustments. Such variability poses a constraint on reproducibility, given that the current parameter tuning process is largely empirical and lacks generalizability across diverse scenarios. Future work will first focus on developing adaptive optimization strategies—specifically designing dynamic adjustment mechanisms for the positive sample mixing ratio based on real-time data quality and noise levels—to enhance the model’s robustness, generalization, and reproducibility across tasks.
@@ Acknowledgement
During the preparation of this work the author(s) used chatgptchatgpt in order to polish our paper. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.